OpenStudy (anonymous):

factor 1 -- 16y ^2 - 2y + 16

OpenStudy (anonymous):

its really simple use quadratic formula

OpenStudy (anonymous):

well i know how to do it but not with the division symbol

OpenStudy (anonymous):

1/(16y^2-2y+16) is it like this???

OpenStudy (mertsj):

Factor a 2 out of the denominator

OpenStudy (anonymous):

is it 4

OpenStudy (anonymous):

the 16 is 4 * 4 = 16

OpenStudy (mertsj):

\[\frac{1}{16y^2-2y+16}=\frac{1}{2(8y^2-y+8)}\]

OpenStudy (anonymous):

thats the answer ?

OpenStudy (mertsj):

Well, it won't factor any further unless you are supposed to find the non-real factors of the denominator. Exactly what do the directions say?

OpenStudy (anonymous):

factor that's all

OpenStudy (mertsj):

Then that's it.